Mr. Jacobs is 4 times as old as his son. Five years ago, he was 9 times as old as his son. Find Mr. Jacob's age.
Answer by w_parminder(53)

Let present age of Jacob's Son = x years
So Jacob's Present age = 4x
5 years ago, Jacob's age = 4x-5
His Son's age = x-5
According to the question,
4x-5 = 9*(x-5)
4x-5 = 9x-45
5x = 40
x = 8
So Jacob's Son's present age = 8 years
And Jacob's present age = 4*8 = 32 years